OpenStudy (yellowlegoguy99):

A spinner is divided into 5 equal segments. The segments are labeled with the letters A, E, M, O, T. Match each compound event with its probability. Round to the nearest percent, if necessary. Terms: P(vowel and consonant) P(M and vowel) P(a letter before J and a letter after R) P(A and A) Definitions: 24% 12% 8% 4%

OpenStudy (yellowlegoguy99):

@tester

OpenStudy (yellowlegoguy99):

@eh

OpenStudy (tester97):

wat?!?!?!?!? O.o i dont remember this stuff ahhhhhhhh wheres the back button?

OpenStudy (yellowlegoguy99):

SOMEONE ESE PLEASE HELP D:

OpenStudy (anonymous):

1) 3/5 * 2/5 = 6/25 = 24% 2) 1/5 * 3/5 = 3/25 =6% 3) 2/5 * 1/5 = 2/25 = 8% 4) 1/5 * 1/5 = 1/25 =4%

OpenStudy (anonymous):

ooooh you should totally give me a medal and be my fan :))) you too @tester97

candycove (candycove):

Where did you get 6% from? That's not even in the options?